OpenStudy (ashleyisakitty):

Help! Halfway there. Two square pyramids have the same volume. For the first pyramid, the side length of the base is 16 in and the height is 28. On the 2nd pyramid the height is 112. What is the side length of the base of the 2nd pyramid?

OpenStudy (ashleyisakitty):

V=1/3(S^2)(h) V=(1/3)(16^2)(28) V=2389 2389=(1/3)(112)(S^2) ??? this is where im starting to get confused.

OpenStudy (ashleyisakitty):

@Compassionate would you know how to do this?

OpenStudy (compassionate):

My account keeps signing me out! really irritating!

OpenStudy (ashleyisakitty):

Hey ive gotten signed out too! It happened only once though.

OpenStudy (compassionate):

It's happened like five times to me. I'm going to write it up in notepad. Okay? Might take awhile because of my connection! Grr

OpenStudy (studygurl14):

Still need help @Ashleyisakitty ?

OpenStudy (ashleyisakitty):

Okay, that sounds good. And yes Studygurl.

OpenStudy (studygurl14):

okay. so hold on a sec while I type it up

OpenStudy (anonymous):

@Ashleyisakitty its basically what i just showed you.

OpenStudy (anonymous):

guys let me have this one?

OpenStudy (anonymous):

so you have 2 equations. 1 to find the volume of both. use the first triangle.

OpenStudy (anonymous):

\[V=1/316^{2}*28\] thats for your volume

OpenStudy (anonymous):

so it comes out to 2389.33

OpenStudy (ashleyisakitty):

Yes.

OpenStudy (anonymous):

Use that as your Volume for the second equation.

OpenStudy (anonymous):

Soooo it would look like this \[2389.33=1/3x ^{2}*112\]

OpenStudy (compassionate):

Yes, so you solve for S: 112/3 = 37.33, so: 2390 = 37.33*(S^2) 64.02 = S^2 S = 8 in

OpenStudy (ashleyisakitty):

so 7.9

OpenStudy (anonymous):

Well did it say to round?

OpenStudy (ashleyisakitty):

Great. awesome. Thank you everyone. Medals for everyone!

OpenStudy (ashleyisakitty):

Yes im going to round up.

OpenStudy (anonymous):

Oh ok welcome . ^-^

OpenStudy (compassionate):

Sorry I was late. It kept signing me out.

OpenStudy (ashleyisakitty):

No worries! Thanks for the effort :)

OpenStudy (anonymous):

@Compassionate i gave you a medal for helping too

OpenStudy (compassionate):

#sweg

OpenStudy (studygurl14):

sorry I left @Ashleyisakitty . I had to eat lunch, then I had computer troubles. sorry!